OpenStudy (anonymous):

A cargo plane left an airport at noon and flew toward New York at the average rate of 400mph. At 2pm a passenger plane left at the same airport for New York and flew the same route as the cargo plane at the average rate of 560mph How many miles did the passenger plane fly before it overtook the cargo plane?

OpenStudy (anonymous):

Help :(

OpenStudy (anonymous):

2800 miles

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

say, the passenger plane flew for "t" hours before it overtook the cargo plane now, the number of miles covered by the cargo plane in that time= 400(t+2), because it had been flying since noon and the distance covered by the passenger plane = 560(t), because it started at 2 pm, now when the passenger plane overtook the cargo plane, it must have been at the same distance from the source airport as was the cargo jet, hence, we equate the distances, i.e, 560t=400(t+2)

OpenStudy (anonymous):

r t = d Let t be the flying time for the passenger plane. The clock for t starts at 2pm, the departure time of day for the passenger plane. At that time the cargo plane has flown 400*2 miles. Solve the following for t, the time to intercept the cargo plane.\[400*t + 400*2=560*t\]\[t=5 \text{ hours}\]The passenger plane flew 560*5 = 2800 miles